OpenStudy (cookielate):

The population of the United States is about 3x10 8th to the power and The population of Mexico is about 1.2x10 to the 8th power how many times larger is the population of the United States than that of Mexico?

OpenStudy (calculusfunctions):

Division is the inverse operation of multiplication. Thus to determine how many greater quantity A is than quantity B, divide quantity A by quantity B.

OpenStudy (cookielate):

so i have to divide 300000000 by 1.200000000

OpenStudy (calculusfunctions):

Correct\[\frac{ 3.0\times10^{8} }{ 1.2\times10^{8} }\]

OpenStudy (cookielate):

idk how to do that

OpenStudy (cookielate):

so i would get 25000000000

OpenStudy (calculusfunctions):

NO, you would just get 2.5

OpenStudy (cookielate):

oh okay

OpenStudy (cookielate):

so how many people live in both the united States and Mexico combined ????

OpenStudy (calculusfunctions):

because 3/(1.2) = 2.5 and (10^8)/(10^8) =1. Thus 2.5 X 1 = 2.5. Understand?

OpenStudy (cookielate):

oh i understand

OpenStudy (calculusfunctions):

To determine the total population of both countries, add the two populations.

OpenStudy (cookielate):

wait so i have to add 1.200000000 to 300000000

OpenStudy (calculusfunctions):

To add two (or more) numbers in scientific notation form, first you have to write both numbers with the same powers of 10. This is already done for you in the question. Now just add the 3.0 and 1.2 and the power of 10 will be the same. So what's 3.0 + 1.2 = ? first

OpenStudy (cookielate):

its 4.2

OpenStudy (calculusfunctions):

Correct! So then the total population is\[(3.0\times10^{8})+(1.2\times10^{8})=4.2\times10^{8}\]Understand @cookielate ??

OpenStudy (cookielate):

yes

OpenStudy (calculusfunctions):

Excellent work!! As a teacher I like students who want to learn rather than just wanting answers because many students on this site just want answers and I don't do that because that's not learning so Good Job!!

OpenStudy (calculusfunctions):

Anything else?

OpenStudy (cookielate):

thanks but then i have to multiply 4.2x10 to the power of 8th or do i have to leave it like that

OpenStudy (calculusfunctions):

No, that's fine unless you're asked to express your final answer in standard form, in which case you would write 420000000. Understand?

OpenStudy (cookielate):

thank you so much i really thank you for helping me so much i dont know how to thank you god bless you and i did not know that you were a teacher o.o

OpenStudy (calculusfunctions):

More than welcome! I appreciate that. Yes I'm a teacher of math/Physics. I log in whenever I can during my spare time to help as many students as possible.

OpenStudy (calculusfunctions):

Can I help you with anything else?

OpenStudy (cookielate):

that so nice your not like other teachers that are mean and rude

OpenStudy (calculusfunctions):

Haha...Thank you.

OpenStudy (cookielate):

no problem ughh i still have to do a little more of math hw

OpenStudy (cookielate):

well i guess i will try my best

OpenStudy (calculusfunctions):

Alright, I'll be logged in for a bit longer so if you need me just ask. Meanwhile I'll see if there's someone else who can use my assistance. Good Luck! Bye for now.

OpenStudy (cookielate):

bye thank you for helping me